# BugRepro: Enhancing Android Bug Reproduction with Domain-Specific Knowledge Integration

### v1

@@ 6 Discussion
The internal threats mainly lie in potential bugs in our implementations. To mitigate this, we leverage the provided artifacts for ReCDroid and AdbGPT to ensure correctness. For BugRepro, two authors independently review and rigorously test the code.
The external threats mainly lie in the generalizability of the dataset.In our experiments, we collect all released bug reports from existing works and all of them are real-world bugs, which can be representative to some degree. However, the effectiveness of BugRepro on a wider range remains to be evaluated in the future. Moreover, the conclusions derived from using a specific LLM (e.g., DeepSeek) may not generalize to other LLMs. To address this, we conduct additional experiments with GPT-4, an advanced LLM, and observe promising and consistent results.

### v2

@@ 6 Threats to Validity
The internal threat mainly lies in potential bugs in our implementations. To mitigate this, we leverage the provided artifacts for ReCDroid and AdbGPT to ensure correctness. For BugRepro, two authors independently review and rigorously test the code.
The external threat mainly lies in the generalizability of the dataset. In our experiments, we collect all released bug reports from existing works and all of them are real-world bugs, which can be representative to some degree. However, the effectiveness of BugRepro on a wider range remains to be evaluated in the future. Moreover, the conclusions derived from using a specific LLM (e.g., DeepSeek) may not generalize to other LLMs. To address this, we conduct additional experiments with GPT-4, an advanced LLM, and observe promising and consistent results.